Given that the amount in an account earning continuous interest is given by the formula A=Pe^(rt) determine the of money that should be invested at a rate of 5% compounded continuously to produce a final balance of $18,000 in 25 years.
Answer by lwsshak3(11628)

Given that the amount in an account earning continuous interest is given by the formula A=Pe^(rt) determine the of money that should be invested at a rate of 5% compounded continuously to produce a final balance of $18,000 in 25 years.
**
P=initial investment
r=annual interest rate=5%
A=amt after t years
t=25 years
..
18000=P*e^(.05*25)
P=18000/e^1.25
using calculator
P=5157
Amt of money that should be invested=$5157